# I see what you mean Co-Speech Gestures for Reference Resolution in Multimodal Dialogue

### v1

@@ 6 Conclusion
In this work, we have studied representational co-speech gestures in collaborative dialogue, using an existing dataset of face-to-face interactions collected by cognitive scientists. We introduced a novel reference resolution task formulated as the problem of identifying the intended referent of a co-speech gesture, while addressing key challenges in gesture representation learning. We proposed a self-supervised Transformer-based approach to learning pre-trained gesture embeddings by jointly exploiting skeletal information and concurrent language encoded with text or speech large language models. Our experiments showed that the resulting gesture embeddings effectively contribute to reference resolution. Representing gestures by exclusively exploiting skeletal information has significant predictive power, and grounding body movements in concurrent speech during pre-training further improves resolution accuracy, even when speech is not provided at test time. Moreover, we showed that reference resolution from representational gestures can benefit from having access to gestures previously used within a dialogue, thus providing empirical support for the presence of gestural entrainment in face-to-face interaction.
Taken together, these findings emphasize the multimodal character of conversation Holler and Levinson (2019b); Özyürek (2014) and the importance of capturing the complementarity between gestures and speech in naturalistic human-machine interaction. Further work is needed to test the extent to which the proposed pre-training approach would transfer to other referential domains and other tasks—a step we leave to future research.
@@
The current work focuses on Dutch-speaking task-oriented dialogues, thus contributing to linguistic diversity in the current English-centric NLP landscape. We nevertheless acknowledge that it is an open question how well the proposed models may generalise to other languages, cultural contexts, tasks, as well as open-domain dialogues. On the methodological front, while we employ and adapt a state-of-the-art motion encoder and show that our pre-training objectives and architecture choices are effective, further optimisation and integration with more advanced speech and semantic encoders may give additional improvements. Finally, the current approach does not ground the reference resolution models in the visual properties of the referents, which might further improve resolution performance. We leave this aspect to future work.
@@ Acknowledgements
We thank the members of the Dialogue Modelling Group at the ILLC, University of Amsterdam, and the Multimodal Language Department at the Max Planck Institute for Psycholinguistics for their feedback. Raquel Fernández acknowledges support from the European Research Council, ERC Consolidator Grant No. 819455.
@@ Appendix A Objects in the CABB Dataset
Figure 9: Distribution of gestures across rounds of interaction.
Appendix B Model Details
B.1 Pre-training Objectives
@@
Table 1: Parameters for the three architectures. The skeleton encoders (the adapted DSTFormer) and the projection heads are the trainable parameters. The speech (i.e., wav2vec2) and text-based semantic (i.e., BERTje) encoders are frozen during pre-training).
@@
Appendix C Implementation Details
@@
C.1 Implementation of the skeleton encoder

### v3

@@ 6 Conclusion
In this work, we have studied representational co-speech gestures in collaborative dialogue, using an existing dataset of face-to-face interactions collected by cognitive scientists. We introduced a novel reference resolution task formulated as the problem of identifying the intended referent of a co-speech gesture, while addressing key challenges in gesture representation learning. We proposed a self-supervised Transformer-based approach to learning pre-trained gesture embeddings by jointly exploiting skeletal information and concurrent language encoded with text or speech large language models. Our experiments showed that the resulting gesture embeddings effectively contribute to reference resolution. Representing gestures by exclusively exploiting skeletal information has significant predictive power, and grounding body movements in concurrent speech during pre-training further improves resolution accuracy, even when speech is not provided at test time. An interesting avenue for future work could be to ground the reference resolution models in the visual properties of the referents, in order to learn mappings between iconic gestures and the objects they represent, which might lead to further improvements. Moreover, we showed that reference resolution from representational gestures can benefit from having access to gestures previously used within a dialogue, thus providing empirical support for the presence of gestural entrainment in face-to-face interaction.
Taken together, our findings emphasize the multimodal character of conversation Holler and Levinson (2019b); Özyürek (2014) and the importance of capturing the complementarity between gestures and speech in naturalistic human-machine interaction. Further work is needed to test the extent to which the proposed pre-training approach would transfer to other referential domains and other tasks—a step we leave to future research.
@@
The current work focuses on Dutch-speaking task-oriented dialogues, thus contributing to linguistic diversity in the current English-centric NLP landscape. We nevertheless acknowledge that it is an open question how well the proposed models may generalise to other languages, cultural contexts, tasks, as well as open-domain dialogues. On the methodological front, while we employ and adapt a state-of-the-art motion encoder and show that our pre-training objectives and architecture choices are effective, further optimisation and integration with more advanced speech and semantic encoders may give additional improvements. Finally, our method is agnostic as to whether concurrent speech and gesture are semantically congruent (i.e., express compatible content). We leverage both information streams and observe that this yields stronger representations and higher performance. Although the nature of the collaborative referential task in the CABB datasets makes it likely that the two modalities align in content, whether the performance gains stem from true cross-modal congruence remains an open empirical question that could be explored in the future. This, however, requires manual annotation of the linguistic referential expressions, which is currently not available.
@@ Acknowledgements
We thank the members of the Dialogue Modelling Group at the ILLC, University of Amsterdam, and the Multimodal Language Department at the Max Planck Institute for Psycholinguistics for their feedback. This project was partially carried out while Esam Ghaleb was at the University of Amsterdam, funded by the Dutch Research Council (NWO) through a Gravitation grant (024.001.006) to the Language in Interaction consortium. Raquel Fernández acknowledges support from the European Research Council, ERC Consolidator Grant No. 819455.
@@ Appendix A Objects in the CABB Dataset
Figure 10: Distribution of manually segmented gestures across rounds of interaction.
Appendix B Segmentation Qualitative Results
Figure 11: Durations distribution of manually and automatically segmented gestures.
Figure 12: Distribution of automatically segmented gestures across rounds of interaction.
Figure 13: Screenshot of ELAN ela (2024), a popular annotation software, which we used to inspect the quality of automatic segmentation.
Figure 11 compares the duration distributions of manually segmented gestures in CABB-S with those automatically segmented in CABB-L. We can see that both distribution curves peak between 0.8 and 0.9 seconds and share a right skew. The automatic segmentation shows a small portion of gestures lasting longer than two seconds. This is most likely because the segmentation model was trained on extended segments to account for the preparation and retraction phases of gestures. Figure 12 plots the average number of gestures per pair across the six rounds of the referential game. Consistent with Figure 10, gesture frequency is highest in the initial rounds and declines in later ones. The higher gesture count in CABB-L reflects longer interaction time: besides the referential interaction, these pairs also carried out an on-screen object localisation task that required identifying the object’s position with respect to the others. We also imported the automatically segmented gestures into ELAN ela (2024) for quality check. As shown in Figure 13, we could visually inspect and verify that gestures’ onsets were detected and segmented with high reliability.
Appendix C Model Details
C.1 Pre-training objectives
@@
C.2 Handling mismatched temporal resolutions.
Multimodal data has different resolutions. Skeletal input and speech raw waveform come with high-frequency motion frames, whereas linguistic information is tokenised into subword units of much coarser granularity. We experiment with two integration strategies.
  1. 1.
    Multimodal (mean-pool). All subword embeddings in an utterance or speech waveform representations of 25 milliseconds are mean-pooled into a single vector. Likewise, the $T$ gesture-frame embeddings are concatenated and projected down to one vector. The two global representations are then fused by the cross-modal Transformer.
  2. 2.
    Multimodal-X (frame-wise attention). Each gesture frame ($\mathbf{q}_{t}$) attends to the set of subword embeddings ($\mathbf{k}_{j},\mathbf{v}_{j}$) for the co-occurring utterance through multi-head attention. We apply the same mechanisms when we handle speech frame-level representations, which operate at segments of 25 milliseconds. This enables more fine-grained alignment, as each frame can focus on the specific linguistic context it co-occurs with.
@@
Table 1: Parameters for the three architectures. The skeleton encoders (the adapted DSTFormer) and the projection heads are the trainable parameters. The speech (i.e., wav2vec2) and text-based semantic (i.e., BERTje) encoders are frozen during pre-training.
Appendix D Implementation Details
@@
D.1 Implementation of the skeleton encoder
@@
Appendix E Impact of Errors on Pose Estimation
Model variant $\sigma$ Accuracy (mean $\pm$ s.d.)
Multimodal-X (clean) 0.00 0.19 $\pm$ 0.05
Multimodal-X (jittered) 0.20 0.18 $\pm$ 0.04
1.00 0.19 $\pm$ 0.05
15.0 0.16 $\pm$ 0.04
Table 2: Reference resolution accuracy using the multimodal-X architecture with skeletal information including different degrees of noise (jittered).
There may be errors in pose estimation because we rely on off-the-shelf ViTPose Xu et al. (2022) and, in the CABB dataset set-up, some joints are occasionally occluded. Several design choices were therefore made to make the models’ representations robust. First, we employ self-supervised objectives, which are known to decrease the effect of noisy inputs Hendrycks et al. (2019). Second, we feed the per-joint confidence scores returned by ViTPose as an additional channel in the skeletal input (see Appendix C), so the model can learn to down-weight unreliable joints.
Additional noise experiment.
To test the models’ resilience to noisy data, we conduct an additional experiment by adding Gaussian noise with varying jitter ($\sigma$) into the 2D skeleton coordinates. Table 2 shows the reference resolution accuracies achieved by the skeleton encoder pre-trained with multimodal-X when different degrees of noise are added. The results show that our model is reasonably robust and only at very high noise levels ($\sigma>10$, which exceeds the typical error rate of current pose estimators) does performance drop slightly.